### v1

@@ 2.7 Estimating local haplotype diversity
More exactly, we perform semi-global sequence-to-DAWG alignment (i.e. requiring the query sequence to be aligned from end to end) by applying BWA-SW to the graph of the linear query sequence $P$. We can find the set of matching haplotypes $\mathcal{M}\triangleq\{v\in V(G_{T}):H_{u_{0}v}>0\}$ where $u_{0}$ represents the start of $P$. $\mathcal{M}$ may include suboptimal haplotypes caused by small variants as well as the optimal one. Importantly, $P$ may be aligned to similar positions on $T$ with slightly different gap placements and alignment scores. In theory, we can identify this case by locating the position of each alignment. This procedure is slow as the “locate” operation is costly. We instead leverage the bi-directionality to identify redundancy heuristically. Suppose $P$ can be aligned to SA bi-intervals $(k_{1},k^{\prime}_{1},s_{1})$ and $(k_{2},k^{\prime}_{2},s_{2})$ and the alignment score to the first interval is higher. We filter out the second interval if $[k_{2},k_{2}+s_{2})\subset[k_{1},k_{1}+s_{1})$ or $[k^{\prime}_{2},k^{\prime}_{2}+s_{2})\subset[k^{\prime}_{1},k^{\prime}_{1}+s_{1})$.
@@ 3 Results
pfp-thres2 51.7 h 51.5 h 788.1 GB
ropebwt3 20.5 h 469.4 h 83.0 GB
metagraph3 16.9 h 314.8 h 251.0 GB
fulgor4 (lossy) 1.2 h 32.7 h 165.2 GB
CommonBacteria ropebwt3 26.5 d 830.3 d 67.3 GB
@@ 3.1 Performance on index construction
We evaluated the performance of BWT construction on 100 haplotype-resolved human assemblies collected in Li et al., (2024). As we included both strands (Section 2.4), each BWT construction algorithm took about 600 billion as input. The average run length is 141.6. grlBWT (commit 5b6d26a; Díaz-Domínguez and Navarro, 2023) is the fastest algorithm (Table 2) at the cost of more than a terabyte of working disk space including decompressed sequences. Ropebwt3 took 21 hours from input sequences in gzip’d FASTA to the final index. It does not use working disk space and can append new sequences to an existing BWT. However, hardcoded for the DNA alphabet, ropebwt3 does not work with more general alphabets. pfp-thresholds (Rossi et al., , 2022) used more memory than the input sequences. It may be impractical with increased sample size.
@@
To test the scalability of ropebwt3 in preparation for larger incoming pangenome datasets, we constructed the BWT of AllTheBacteria (Hunt et al., , 2024) excluding the “dustbin” and “unknown” categories that include relatively unique genomes. This subset, which we call as CommonBacteria, consists of 278.4 million sequences in 7.33 Tb. An older version of ropebwt3 took less than a month to construct the double-strand BWT with 15.66 trillion symbols. We expect the latest version to reduce the elapsed time to about 17 days. The final index in the fermi format is 27.6 GB in size. Indexing the entire AllTheBacteria dataset will probably take 100–200 GB as unique sequences are not compressed well.
@@ 3.2 Query performance
Among the three BWT-based tools, ropebwt3 is slower than Movi but faster than MONI on finding exact matches. Movi finds pseudo-matching length (PML). Although the fastest, it finds the longest exact matches for only 195 out of one million short reads and it misses the longest exact matches for all long reads. The longest PML of each read is on average 27% shorter than the longest exact match for LR$+$ and 22% shorter for SR$+$. Nonetheless, we believe it should be possible to implement SMEM finding based on the Movi data structure with a minor performance overhead.
@@ 3.3 Identifying novel sequences
As a biological application, we used the pangenome index to identify novel sequences in reads that are absent from other human genomes. For this, we downloaded the PacBio HiFi reads for tumor sample COLO829 (https://downloads.pacbcloud.com/public/revio/2023Q2/COLO829/COLO829/), mapped them to the pangenome index with 100 human genomes and extracted $\geq$1kb regions on reads that are not covered by SMEMs of 51bp or longer. We found 95kb sequences in 43 reads. These sequences could not be assembled. NCBI BLAST suggested multiple weak hits to cow genomes. We could not identify the source of these sequences but there were few of them anyway.
When we mapped the COLO829 reads to a single T2T-CHM13 genome (Nurk et al., , 2022) and applied the same procedure, we found 55.9Mb of “novel” sequences in 25,365 reads. The much larger number is probably caused by regions with dense SNPs that prevented long exact matches. Counterintuitively, mapping these reads to T2T-CHM13 with minimap2 (Li, , 2018) resulted in more “novel” sequences at 78.6Mb, probably because minimap2 ignores seeds occurring thousands of times in T2T-CHM13 and may miss more alignments. Capable of finding SMEMs at the pangenome scale, ropebwt3 is more effective for identifying known sequences. It is also faster than full minimap2 alignment against a single genome.
@@ 3.4 Haplotype diversity around C4 genes
Figure 3: Haplotype diversity around the C4A gene. 101-mers with 50bp overlaps are extracted from the C4A genomic sequence on GRCh38 and aligned to the human pangenome. The Y axis shows the count of a 101-mer matches under different edit-distance thresholds.
@@ 4 Discussions
Ropebwt3 is a fast tool for BWT construction and sequence search for redundant DNA sequences. Generating BWT purely in memory and supporting incremental build, ropebwt3 is convenient and practical for BWT construction at large scale. It provides the fastest algorithm so far for finding supermaximal exact matches and can report inexact hits as well. Ropebwt3 demonstrates that BWT-based data structures are scalable to terabases of pangenome data.
Ropebwt3 will take 2–3 weeks to index 1,000 human genomes assembled from long reads. When there are more genomes, it will be faster to construct the BWT of individual genome independently and then merge them one by one. This strategy will use working disk space but achieve better parallelization. Sirén, (2016) introduced disk-based algorithms for merging large BWTs, which may further improve the parallelization of BWT merge.
We thought about implementing r-index in ropebwt3. However, although an r-index is faster than an FM-index of the same size, it imposes a fixed sampling rate: two suffix array values per run. The BWT of CommonBacteria has 14.6 trillion bases and 17.6 billion runs. An r-index is likely to take more than 200GB, while an FM-index sampled at one position per 8,192bp takes 17.5GB in ropebwt3. We are still looking for a more balanced algorithm for the “locate” operation.
We often use pangenome graphs to analyze multiple similar genomes (Liao et al., , 2023). These graphs are built from the multiple sequence alignment through complex procedures involving many parameters (Li et al., , 2020; Hickey et al., , 2023; Garrison et al., , 2023). It is challenging to understand if the graph topology is biologically meaningful especially given that we do not know the correct alignment between two genomes, let alone multiple ones. Complement to graph-based data structures, BWT-based algorithms are often exact with no heuristics or parameters but they tend to support limited query types. For example, we cannot project the alignment to a designated reference genome. What additional query types we can achieve will be of great interest to the comprehensive pangenome analysis in future.
@@ Acknowledgments
The ropebwt3 source code is available at https://github.com/lh3/ropebwt3. Prebuilt ropebwt3 indices can be obtained from https://zenodo.org/doi/10.5281/zenodo.11533210.

### v2

@@ 2.8 Estimating local haplotype diversity
More exactly, we perform semi-global sequence-to-DAWG alignment (i.e. requiring the query sequence to be aligned from end to end) by applying BWA-SW to the graph representing the linear query sequence $P$, from the end to the start. We can find the set of matching haplotypes $\mathcal{M}\triangleq\{v\in V(G_{T}):H_{u_{0}v}>0\}$ where $u_{0}$ represents the start of $P$. $\mathcal{M}$ may include suboptimal haplotypes caused by small variants as well as the optimal one.
Importantly, $P$ may be aligned to similar positions on $T$ with slightly different gap placements. For example, given $T={\tt CAAGCAG}$ and $P={\tt AGCG}$, the algorithm above may find the following two alignments around the same position but in different SA intervals:
   T: CAAGCAG        CAAGCAG
        ||||    or    | |||
   P:   AGCA          A-GCA
When counting hits, we want to ignore the second suboptimal alignment. In theory, we can identify this case by comparing the position of each alignment. This procedure is slow as the locate operation is costly. We instead leverage the bi-directionality to identify redundancy heuristically. Suppose $P$ can be aligned to SA bi-intervals $(k_{1},k^{\prime}_{1},s_{1})$ and $(k_{2},k^{\prime}_{2},s_{2})$ and the alignment score to the first interval is higher. We filter out the second interval if $[k_{2},k_{2}+s_{2})\subset[k_{1},k_{1}+s_{1})$ or $[k^{\prime}_{2},k^{\prime}_{2}+s_{2})\subset[k^{\prime}_{1},k^{\prime}_{1}+s_{1})$. This strategy does not avoid all overcounting but it works well on multiple real examples we have closely inspected.
@@ 3 Results
Table 2: Datasets
[…]
human3204 963.0 Gb 27.1 k 395.6
CommonBacteria5 7326.6 Gb 278.4 M 828.6
[…]
{tablenotes}
number of bases in the input sequences on one strand
average run length in BWT constructed from both strands
100 long-read human assemblies; N50 string length: 44.4 Mb
320 long-read human assemblies; N50 string length: 135.3 Mb
AllTheBacteria (Hunt et al., , 2024) excluding “dustbin” and “unknown”
@@
pfp-thres3 $<$51.7 h ×$\times 1.0$ 788.1 GB
ropebwt3 20.5 h ×$\times 22.9$ 83.0 GB
metagraph4 16.9 h ×$\times 18.6$ 251.0 GB
fulgor5 1.2 h ×$\times 27.2$ 165.2 GB
human320 grlBWT6 23.3 h ×$\times 4.2$ 270.4 GB
ropebwt3 81.2 h ×$\times 16.5$ 99.2 GB
ropebwt37 64.9 h ×$\times 23.7$ 170.5 GB
CommonBacteria ropebwt3 25.6 d ×$\times 32.4$ 67.3 GB
@@
using two bytes per run with option ‘-b 2’, which is faster
BWT merge and partial BWT construction with libsais are run together
@@ 3.1 Performance on index construction
We evaluated the performance of BWT construction on 100 haplotype-resolved human assemblies collected in Li et al., (2024). As we included both strands (Section 2.5), each BWT construction algorithm took about 600 billion bases as input (Table 2). grlBWT (commit 5b6d26a; Díaz-Domínguez and Navarro, 2023) is the fastest algorithm (Table 3) at the cost of $\sim$2 terabytes of working disk space including decompressed sequences. Ropebwt3 took 21 hours from input sequences in gzip’d FASTA to the final index, of which 7.7 hours was spent on libsais. It does not use working disk space and can append new sequences to an existing BWT. However, hardcoded for the DNA alphabet, ropebwt3 does not work with more general alphabets. pfp-thresholds (Rossi et al., , 2022) used more memory than the input sequences. It may be impractical with increased sample size.
@@
To test scalability, we indexed a larger dataset consisting of 320 human genomes recently released by the Human Pangenome Reference Consortium (Liao et al., , 2023). Because these assemblies are more contiguous than human100 (Table 2), $m_{2}$ in Algorithm 1 is smaller, which reduces the multi-threading efficiency of ropebwt3. We can alleviate this by executing BWT merge and partial BWT construction at the same time at the cost of higher memory footprint. On human320, grlBWT is 2–4 times as fast but uses more memory (Table 3). On the largest CommonBacteria dataset (Table 2), ropebwt3 took less than a month to construct the double-strand BWT with 14.66 trillion symbols. The final index in the fermi format is 27.6 GB in size. The “dustbin” and “unknown” categories in AllTheBacteria consist of many unique sequences which are not compressed well. Indexing the entire AllTheBacteria with ropebwt3 will probably take 100–200 GB memory at the peak.
@@ 3.2 Query performance
Among the three BWT-based tools, ropebwt3 is slower than Movi but faster than MONI on finding exact matches. Movi finds pseudo-matching length (PML) which is not intended to be the longest exact match. PML corresponds to the longest MEM for only 195 out of one million short reads, and none for long reads. The longest PML of each read is on average 27% shorter than the longest exact match for LR$+$ and 22% shorter for SR$+$. Nonetheless, we believe it is possible to implement SMEM finding based on the Movi data structure with minor performance overhead.
@@ 3.3 Identifying novel sequences
As a biological application, we used the pangenome index to identify novel sequences in reads that are absent from other human genomes. For this, we downloaded the PacBio HiFi reads for tumor sample COLO829 (https://downloads.pacbcloud.com/public/revio/2023Q2/COLO829/COLO829/), mapped them to human100 (Table 2), which includes T2T-CHM13 (Nurk et al., , 2022), and extracted $\geq$1kb regions on reads that are not covered by SMEMs of 51bp or longer. We found 95kb sequences in 43 reads. These sequences could not be assembled. NCBI BLAST suggested multiple weak hits to cow genomes. We could not identify the source of these sequences but there were few of them anyway.
When we mapped the COLO829 reads to T2T-CHM13 only and applied the same procedure, we found 55.9Mb of “novel” sequences in 25,365 reads. The much larger number is caused by regions with dense SNPs that prevented long exact matches. Counterintuitively, mapping these reads to T2T-CHM13 with minimap2 (Li, , 2018) resulted in more “novel” sequences at 78.6Mb, probably because minimap2 ignores seeds occurring thousands of times in T2T-CHM13 and may miss more alignments. Capable of finding SMEMs at the pangenome scale, ropebwt3 is more effective for identifying known sequences. It is also 16% faster and uses less memory than full minimap2 alignment against a single genome.
@@ 3.4 Haplotype diversity around C4 genes
Figure 3: Haplotype diversity around the C4A gene. 101-mers with 50bp overlaps are extracted from the C4A genomic sequence on GRCh38 and aligned to the human pangenome. The Y axis shows the counts of 101-mer matches under different edit-distance thresholds.
@@ 4 Discussions
Ropebwt3 is a fast tool for BWT construction and sequence search for redundant DNA sequences. Generating BWT purely in memory and supporting incremental build, ropebwt3 is convenient and practical for BWT construction at large scale. It is the only algorithm that can construct the BWT of 320 human genomes from a 2.5 GB AGC archive (Deorowicz et al., , 2023) without staging all decompressed data in memory or on disk. It provides the fastest algorithm so far for finding supermaximal exact matches and can report inexact hits as well. Ropebwt3 demonstrates that BWT-based data structures are scalable to terabases of pangenome data.
Ropebwt3 implements an FM-index to locate SMEMs or local hits. Although the standard r-index is faster than an FM-index of the same size, it imposes a fixed sampling rate: two suffix array values per run. The BWT of CommonBacteria has 14.6 trillion bases and 17.6 billion runs. An r-index is likely to take more than 200GB, while an FM-index sampled at one position per 8,192bp takes 17.5GB in ropebwt3. Subsampled r-index (sr-index; Cobas et al., 2024) is probably the better solution, which we will explore in future.
This article has not evaluated several recent BWT-based tools including r-index-f (Brown et al., , 2022), block_RLBWT (Díaz-Domínguez et al., , 2023), Move-r (Bertram et al., , 2024) and b-move (Depuydt et al., , 2024). Although they have not been tested on large datasets like human pangenome and they do not report SMEMs, some of their data structures are probably more efficient than the ones we use. We may integrate these methods into ropebwt3 in future as well.
We often use pangenome graphs to analyze multiple similar genomes (Liao et al., , 2023). These graphs are built from the multiple sequence alignment through complex procedures involving many parameters (Li et al., , 2020; Hickey et al., , 2023; Garrison et al., , 2024). It is challenging to understand if the graph topology is biologically meaningful especially given that we do not know the correct alignment between two genomes, let alone multiple ones. Complement to graph-based data structures, BWT-based algorithms are often exact with no heuristics or parameters but they tend to support limited query types. For example, we cannot project the alignment to a designated reference genome. What additional query types we can achieve will be of great interest to the comprehensive pangenome analysis in future.
@@ Acknowledgments
The ropebwt3 source code is available at https://github.com/lh3/ropebwt3. Prebuilt ropebwt3 indices can be obtained from https://doi.org/10.5281/zenodo.11533210 and https://doi.org/10.5281/zenodo.13955431.